OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 10, and minor axis of length 8

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

\[\large \frac{ x^2 }{a^2 }+\frac{ y^2 }{ b^2 }=1\] you're given the vertical axis, and vertical axis = 2b you're given the horizontal axis, and horizontal axis = 2a

OpenStudy (anonymous):

16 and 25?

OpenStudy (agent0smith):

yes, for a^2 and b^2

OpenStudy (anonymous):

x^2/16?

OpenStudy (anonymous):

?

OpenStudy (agent0smith):

\[\large \frac{ x^2 }{a^2 }+\frac{ y^2 }{ b^2 }=1\]

OpenStudy (anonymous):

so im worng

OpenStudy (anonymous):

x^2/25

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

x^2/4+y^2/5

OpenStudy (agent0smith):

you already had a^2 and b^2 way back here... 16 and 25? 3 minutes agoReport Abuse 99 agent0smith Medals 1 yes, for a^2 and b^2